# Symmetric Iterations with Countable Support: A Framework for Choiceless ZF Extensions

### v1

@@ 3.1 Framework and Setup
  • Automorphisms: $\mathcal{G}_{\lambda}$ is the inverse limit of $\langle\mathcal{G}_{\beta},\pi^{\gamma}_{\beta}:\gamma\leq\beta<\lambda\rangle$:
    $\mathcal{G}_{\lambda}\ :=\ \Bigl\{\langle g_{\beta}:\beta<\lambda\rangle:\forall\gamma<\beta<\lambda\ \bigl(\pi^{\gamma}_{\beta}(g_{\beta})=g_{\gamma}\bigr)\Bigr\}.$
    It acts coordinatewise on conditions by
    $(g\cdot p)(\beta)\ :=\ g_{\beta}\cdot p(\beta)\qquad(\beta\in\operatorname{dom}(p)),$
    so in particular $\mathrm{supp}(g\cdot p)=\mathrm{supp}(p)$ and hence supports remain countable.
  • Filter: the limit filter is denoted $\tilde{\mathcal{F}}_{\lambda}$. It is defined in Definition 3.11 and constructed in Section 3.2; the key property is that $\tilde{\mathcal{F}}_{\lambda}$ is normal and $\omega_{1}$–complete.
Stage $\alpha=\lambda$ (Limit, $\mathrm{cf}(\lambda)\geq\omega_{1}$):
For uncountable cofinality, use the direct-limit presentation (Remark 2.7):
  • Forcing: $\mathbb{P}_{\lambda}=\bigcup_{\beta<\lambda}\mathbb{P}_{\beta}$, with canonical projections $\pi^{\beta}_{\lambda}:\mathbb{P}_{\lambda}\to\mathbb{P}_{\beta}$ (restriction to stage $\beta$).
  • Automorphisms: $\mathcal{G}_{\lambda}=\bigcup_{\beta<\lambda}\mathcal{G}_{\beta}$, with canonical restrictions $\pi^{\beta}_{\lambda}:\mathcal{G}_{\lambda}\to\mathcal{G}_{\beta}$.
  • Filter: the limit filter is denoted $\tilde{\mathcal{F}}_{\lambda}$ and is the normal filter generated by head pushforwards:
    $\mathcal{B}_{\lambda}\ :=\ \bigcup_{\beta<\lambda}\pi^{\beta}_{\lambda,*}(\mathcal{F}_{\beta}),\qquad\pi^{\beta}_{\lambda,*}(H)\ :=\ \{g\in\mathcal{G}_{\lambda}:\pi^{\beta}_{\lambda}(g)\in H\}.$
    Equivalently, this is the $\mathrm{cf}(\lambda)\geq\omega_{1}$ clause of Definition 3.11.
In this case, $\tilde{\mathcal{F}}_{\lambda}$ is $\omega_{1}$–complete by stage-bounding (Remark 3.3).
Remark 3.2 (Relation to Karagila’s finite-support framework).
This definition addresses Karagila’s open problem [5] about extending symmetric iterations to countable support. Karagila’s Def. 5.1 in [1] formalizes finite-support symmetric iterations and the productive framework (Def. 8.1) that ensures compositional fidelity through block-by-block tracking via excellent supports. Our Def. 3.1 uses $\omega_{1}$-complete limit filters (Theorem 3.15)—via block factorization at countable cofinality and stage-bounding at uncountable cofinality—to preserve ZF and DC with countable supports. The mechanisms are fundamentally different: Karagila’s productive conditions ensure compositional correctness of finite iterations, while our filter completeness provides the closure properties needed for DC-preservation with infinite supports.
@@
Remark 3.3 (Direct limits preserve $\omega_{1}$-completeness at uncountable cofinality).
Assume $\mathrm{cf}(\lambda)\geq\omega_{1}$. Let $\tilde{\mathcal{F}}_{\lambda}$ be the filter on $\mathcal{G}_{\lambda}$ generated by $\bigcup_{\beta<\lambda}\pi^{\beta}_{\lambda,*}(\mathcal{F}_{\beta})$ in the direct-limit case, where $\mathcal{G}_{\lambda}=\bigcup_{\beta<\lambda}\mathcal{G}_{\beta}$ and pushforwards are inclusions. Then $\tilde{\mathcal{F}}_{\lambda}$ is $\omega_{1}$-complete.
Indeed, given a countable family $\{H_{n}:n<\omega\}\subseteq\tilde{\mathcal{F}}_{\lambda}$, choose $\beta_{n}<\lambda$ and $K_{n}\in\mathcal{F}_{\beta_{n}}$ with $H_{n}=\pi^{\beta_{n}}_{\lambda,*}(K_{n})=K_{n}$ (the equality since, in a direct limit, pushforwards are inclusions). The set $\{\beta_{n}:n<\omega\}$ is countable and hence bounded in $\lambda$ because $\mathrm{cf}(\lambda)\geq\omega_{1}>\omega$; equivalently, if it were unbounded then $\mathrm{cf}(\lambda)\leq\omega$, a contradiction. Thus $\beta^{*}:=\sup_{n}\beta_{n}<\lambda$.
Since $\beta_{n}\leq\beta^{*}$ for each $n$, we have $K_{n}\in\mathcal{F}_{\beta_{n}}\subseteq\mathcal{F}_{\beta^{*}}$. By the induction hypothesis, $\mathcal{F}_{\beta^{*}}$ is $\omega_{1}$-complete, so $\bigcap_{n<\omega}K_{n}\in\mathcal{F}_{\beta^{*}}\subseteq\tilde{\mathcal{F}}_{\lambda}$. Therefore $\bigcap_{n<\omega}H_{n}=\bigcap_{n<\omega}K_{n}\in\tilde{\mathcal{F}}_{\lambda}$.
Remark 3.4 (Key Properties to Verify).
For the iteration to be well-defined, we must verify at each stage:
  1. 1.
    Automorphisms are well-defined
  2. 2.
    Filter is proper (non-empty)
  3. 3.
    Filter is normal (closed under conjugation)
  4. 4.
    Filter is $\omega_{1}$-complete (limit stages)
Items 1–3 follow from the definitions and Theorem 3.15; Item 4 is the technical heart of the paper.
Remark 3.5 (Comparison to Karagila’s Finite Support).
In Karagila’s finite support framework [1], supports are finite and filters need only be closed under finite intersections (automatic). Countable support requires $\omega_{1}$-completeness, which is nontrivial and the core contribution of this framework.
Example 3.6 (Iteration Setup).
For a Cohen iteration to $\omega_{1}$:
  • $\dot{\mathbb{Q}}_{\beta}=\mathrm{Add}(\omega,2)$
  • Automorphisms: finite-support permutations of Cohen coordinates
  • Filter: cofinite stabilizers (e.g., fix all but finitely many coordinates)
@@ 3.2 Filter Generation at Limit Stages
Let $\lambda$ be a limit ordinal with $\mathrm{cf}(\lambda)=\omega$, and assume each $\mathcal{F}_{\beta}$ for $\beta<\lambda$ is a normal, $\omega_{1}$-complete filter on $\mathcal{G}_{\beta}$ (by induction). We construct $\mathcal{F}_{\lambda}$ so that:
  • Extends predecessors: Each $\mathcal{F}_{\beta}$ embeds into $\mathcal{F}_{\lambda}$ via the natural projection.
  • Normal: $\mathcal{F}_{\lambda}$ is closed under conjugation by $\mathcal{G}_{\lambda}$.
  • $\omega_{1}$-complete: $\mathcal{F}_{\lambda}$ is closed under countable intersections.
  • Generated: $\tilde{\mathcal{F}}_{\lambda}$ is generated by head pushforwards from earlier stages; at $\mathrm{cf}(\lambda)=\omega$ we additionally close under countable intersections in the minimal normal way.
Definition 3.7 (Pushforward Filter).
For $\beta<\lambda$, let $\pi^{\beta}_{\lambda}:\mathcal{G}_{\lambda}\to\mathcal{G}_{\beta}$ be the projection. Define the *pushforward filter* $\pi^{\beta}_{\lambda,*}(\mathcal{F}_{\beta})$ on $\mathcal{G}_{\lambda}$ to be the filter generated by
$\bigl\{\{g\in\mathcal{G}_{\lambda}:\pi^{\beta}_{\lambda}(g)\in H\}:H\in\mathcal{F}_{\beta}\bigr\}.$
Equivalently, each $H\in\mathcal{F}_{\beta}$ yields the generator $\{g:\pi^{\beta}_{\lambda}(g)\in H\}$ in $\mathcal{G}_{\lambda}$.
Remark 3.8.
The pushforward $\pi^{\beta}_{\lambda,*}(\mathcal{F}_{\beta})$ is sometimes called a head pushforward, since it lifts symmetry from an earlier stage $\beta$ to a later stage $\lambda$.
@@
Definition 3.11 (Filter at a limit stage).
Let $\lambda$ be a limit ordinal and assume we have already defined the symmetric systems $(\mathbb{P}_{\beta},\mathcal{G}_{\beta},\mathcal{F}_{\beta})$ for all $\beta<\lambda$. Let
$\mathcal{B}_{\lambda}\ :=\ \bigcup_{\beta<\lambda}\pi^{\beta}_{\lambda,*}(\mathcal{F}_{\beta})\qquad\text{where}\qquad\pi^{\beta}_{\lambda,*}(H)=\{g\in\mathcal{G}_{\lambda}:\pi^{\beta}_{\lambda}(g)\in H\}.$
@@
    If $\mathrm{cf}(\lambda)\geq\omega_{1}$, define $\tilde{\mathcal{F}}_{\lambda}$ to be the normal filter generated by $\mathcal{B}_{\lambda}$, i.e.
@@
    If $\mathrm{cf}(\lambda)=\omega$, define $\tilde{\mathcal{F}}_{\lambda}$ to be the normal $\omega_{1}$–complete filter generated by $\mathcal{B}_{\lambda}$, i.e.
@@
Remark. The second clause is the minimal normal way to add $\omega_{1}$–completeness at $\mathrm{cf}(\lambda)=\omega$ limits, without introducing extra generators tied to arbitrary name-stabilizers.
@@
The $\omega_{1}$–completeness argument splits by cofinality. If $\mathrm{cf}(\lambda)\geq\omega_{1}$, stage bounding reduces countable intersections to a single stage filter. If $\mathrm{cf}(\lambda)=\omega$, stage bounding fails and we instead use tuple–stabilizer compression at the inverse limit.
The proof of $\omega_{1}$-completeness requires different arguments depending on the cofinality of $\lambda$:
  • Case $\mathrm{cf}(\lambda)\geq\omega_{1}$ (direct limit). Here stage-bounding applies: any countable set of ordinals below $\lambda$ is bounded, since a countable cofinal subset would contradict $\mathrm{cf}(\lambda)\geq\omega_{1}$. This case is handled in Remark 3.3.
  • Case $\mathrm{cf}(\lambda)=\omega$ (inverse limit). Stage-bounding fails: a countable set can be cofinal in $\lambda$. We instead use block factorization along a fixed cofinal sequence.
@@
When $\mathrm{cf}(\lambda)=\omega$, the stage-bounding lemma fails. For example, if $\lambda=\omega$, the set $\{n:n<\omega\}$ is countable but $\sup_{n}n=\omega=\lambda$. Similarly, if $\lambda=\aleph_{\omega}$, the sequence $\{\aleph_{n}:n<\omega\}$ is countable and cofinal. This is precisely why we need stabilizer compression rather than stage-bounding at limits of countable cofinality.
Theorem 3.15 (Filter Extension Theorem).
Let $\lambda$ be a limit stage and let $\tilde{\mathcal{F}}_{\lambda}$ be defined as in Definition 3.11. Assume inductively that each $\mathcal{F}_{\beta}$ for $\beta<\lambda$ is a normal $\omega_{1}$–complete filter. Then:
@@
    $\tilde{\mathcal{F}}_{\lambda}$ is a normal filter on $\mathcal{G}_{\lambda}$ extending each head pushforward $\pi^{\beta}_{\lambda,*}(\mathcal{F}_{\beta})$.
@@
Normality and extension are immediate from the generating descriptions in Definition 3.11.
For $\omega_{1}$–completeness, let $\{H_{n}:n<\omega\}\subseteq\tilde{\mathcal{F}}_{\lambda}$.
Case $\mathrm{cf}(\lambda)=\omega$. For each $n$ choose sequences $\langle B^{n}_{m}:m<\omega\rangle\in\mathcal{B}_{\lambda}^{\omega}$ and $\langle g^{n}_{m}:m<\omega\rangle\in\mathcal{G}_{\lambda}^{\omega}$ witnessing $H_{n}$, i.e.
$\bigcap_{m<\omega}g^{n}_{m}B^{n}_{m}(g^{n}_{m})^{-1}\ \leq\ H_{n}.$
Enumerate $\omega\times\omega$ as a single $\omega$–sequence and take the corresponding countable family of conjugates. Their intersection is contained in $\bigcap_{n<\omega}H_{n}$, hence $\bigcap_{n<\omega}H_{n}\in\tilde{\mathcal{F}}_{\lambda}$ by Definition 3.11.
Case $\mathrm{cf}(\lambda)\geq\omega_{1}$. Each $H_{n}$ contains a finite intersection of conjugates of members of $\mathcal{B}_{\lambda}$, hence depends on only finitely many stages. Let $\beta_{n}<\lambda$ bound the stages appearing in a witness for $H_{n}$; set $\beta^{*}=\sup_{n<\omega}\beta_{n}<\lambda$ (using $\mathrm{cf}(\lambda)\geq\omega_{1}$). Transport each witness to stage $\beta^{*}$ using the coherence of the projections, apply $\omega_{1}$–completeness of $\mathcal{F}_{\beta^{*}}$, and pull back to $\lambda$. This yields $\bigcap_{n<\omega}H_{n}\in\tilde{\mathcal{F}}_{\lambda}$. ∎

### v3

@@ 3.1 Framework and Setup
$\mathcal{G}_{\lambda}$ is the inverse limit of $\langle\mathcal{G}_{\beta},\rho_{\gamma,\beta}:\gamma<\beta<\lambda\rangle$ and acts coordinatewise on $\mathbb{P}_{\lambda}$ (Lemma 3.3). The limit filter $\tilde{\mathcal{F}}_{\lambda}$ is defined in Definition 3.9.
Limit stage $\lambda\leq\Theta$ with $\mathrm{cf}(\lambda)\geq\omega_{1}$.
We identify the countable-support limit with the direct limit (Remark 3.2):
$\mathbb{P}_{\lambda}=\bigcup_{\beta<\lambda}\mathbb{P}_{\beta},\qquad\mathcal{G}_{\lambda}=\bigcup_{\beta<\lambda}\mathcal{G}_{\beta},$
with the canonical restriction maps $\pi_{\beta,\lambda}(p)=p\upharpoonright\beta$ and $\rho_{\beta,\lambda}$. The limit filter $\tilde{\mathcal{F}}_{\lambda}$ is the normal filter generated by head pullbacks from earlier stages (Definition 3.9), and it is $\omega_{1}$-complete by stage-bounding (Remark 3.5).
Remark 3.2 (Bookkeeping at uncountable cofinality).
Let $\lambda$ be a limit ordinal with $\mathrm{cf}(\lambda)\geq\omega_{1}$, and let $\mathbb{P}_{\lambda}$ be the usual countable-support limit of $\langle\mathbb{P}_{\alpha},\dot{\mathbb{Q}}_{\alpha}:\alpha<\lambda\rangle$. Then every $p\in\mathbb{P}_{\lambda}$ has bounded support: since $\mathrm{supp}(p)$ is countable and $\mathrm{cf}(\lambda)>\omega$, there is $\beta<\lambda$ with $\mathrm{supp}(p)\subseteq\beta$, hence $p\in\mathbb{P}_{\beta}$. Therefore $\mathbb{P}_{\lambda}=\bigcup_{\beta<\lambda}\mathbb{P}_{\beta}$ (direct limit identification), and similarly $\mathcal{G}_{\lambda}=\bigcup_{\beta<\lambda}\mathcal{G}_{\beta}$ with the evident restriction maps.
Lemma 3.3 (Limit-stage coordinatewise action).
Let $\lambda$ be a limit stage with $\mathrm{cf}(\lambda)=\omega$. If $g=\langle g_{\beta}:\beta<\lambda\rangle\in\mathcal{G}_{\lambda}$, define $(g\cdot p)(\beta)=g_{\beta}\cdot p(\beta)$ for $\beta\in\operatorname{dom}(p)$. Then $p\mapsto g\cdot p$ is an automorphism of $\mathbb{P}_{\lambda}$.
Proof.
It is immediate that $\operatorname{dom}(g\cdot p)=\operatorname{dom}(p)$ and $g\cdot 1_{\mathbb{P}_{\lambda}}=1_{\mathbb{P}_{\lambda}}$. For order preservation, suppose $q\leq p$ in $\mathbb{P}_{\lambda}$. Fix $\beta\in\operatorname{dom}(p)$. Then $q\upharpoonright\beta\Vdash_{\mathbb{P}_{\beta}}q(\beta)\leq p(\beta)$. Apply the invariance of the forcing relation under automorphisms from Section 2 (Lemma 2.9 in its forcing-theorem form): since $g\upharpoonright\beta\in\mathcal{G}_{\beta}$,
$(g\cdot q)\upharpoonright\beta\ \Vdash_{\mathbb{P}_{\beta}}\ (g\cdot q)(\beta)\leq(g\cdot p)(\beta).$
As this holds for all $\beta\in\operatorname{dom}(p)$, we have $g\cdot q\leq g\cdot p$. Bijectivity follows from applying the same argument to $g^{-1}$. ∎
Remark 3.4 (Comparison with related frameworks).
The successor-step symmetric-system construction used above is standard (see, e.g., [3, §§3–5]). What is new in the present paper is the limit-stage treatment needed for countable support: at $\mathrm{cf}(\lambda)\geq\omega_{1}$ the direct-limit identification (Remark 3.2) allows stage-bounding, while at $\mathrm{cf}(\lambda)=\omega$ we define the limit filter by taking the smallest normal $\omega_{1}$-complete extension of the head pullbacks (Definition 3.9).
The independent concurrent work of Karagila–Schilhan [6] develops a broad structural theory of symmetric extensions—including products, reduced products, quotients, and a general $I$-support iteration framework (where $I$ is any ideal on the index ordinal)—and proves results on Kinna–Wagner principles and the HOD theorem. Their $I$-support definition subsumes countable support as a special case. The present paper is complementary in focus: we prove that the specific $\omega_{1}$-complete normal limit filter construction at cofinality-$\omega$ stages (Theorem 3.13) yields DC preservation (Theorem 3.20), which is not addressed in [6]. Accordingly, the present paper should be read as a limit-stage supplement to the standard successor-step theory, rather than as a replacement for more general support frameworks.
@@
Remark 3.5 (Uncountable cofinality: stage-bounding yields $\omega_{1}$-completeness).
Assume $\mathrm{cf}(\lambda)\geq\omega_{1}$ and work in the direct-limit identification of Remark 3.2. Let $\tilde{\mathcal{F}}_{\lambda}$ be the limit filter from Definition 3.9(1), i.e. the smallest normal filter on $\mathcal{G}_{\lambda}$ containing the head pullbacks $\rho_{\beta,\lambda}^{-1}(H)$ for $\beta<\lambda$ and $H\in\mathcal{F}_{\beta}$.
Then $\tilde{\mathcal{F}}_{\lambda}$ is $\omega_{1}$-complete. Indeed, let $\langle K_{n}:n<\omega\rangle\subseteq\tilde{\mathcal{F}}_{\lambda}$. For each $n$, by the normal-generation clause in Definition 3.9(1), fix witnesses $k_{n}<\omega$, ordinals $\beta_{n,i}<\lambda$, groups $H_{n,i}\in\mathcal{F}_{\beta_{n,i}}$, and elements $g_{n,i}\in\mathcal{G}_{\lambda}$ such that
$\bigcap_{i<k_{n}}g_{n,i}\,\rho_{\beta_{n,i},\lambda}^{-1}(H_{n,i})\,g_{n,i}^{-1}\ \leq\ K_{n}.$
Since $\mathrm{cf}(\lambda)\geq\omega_{1}$, the set $\{\beta_{n,i}:n<\omega,\ i<k_{n}\}$ is countable and hence bounded in $\lambda$; let $\beta^{*}<\lambda$ bound it.
For each $n,i$, set $H^{\prime}_{n,i}:=\rho_{\beta_{n,i},\beta^{*}}^{-1}(H_{n,i})$. Then $H^{\prime}_{n,i}\in\mathcal{F}_{\beta^{*}}$ by Lemma 3.7, and
$\rho_{\beta^{*},\lambda}^{-1}(H^{\prime}_{n,i})\ \leq\ \rho_{\beta_{n,i},\lambda}^{-1}(H_{n,i})$
by coherence of restrictions. Replacing each head pullback by $\rho_{\beta^{*},\lambda}^{-1}(H^{\prime}_{n,i})$ only strengthens the left-hand side, so we may assume all generators come from stage $\beta^{*}$.
Now use $\omega_{1}$-completeness of $\mathcal{F}_{\beta^{*}}$ (by induction on stages, using the successor-stage $\omega_{1}$-completion) to form the countable intersection of the corresponding (conjugated) stage-$\beta^{*}$ generators inside $\mathcal{F}_{\beta^{*}}$, and pull back along $\rho_{\beta^{*},\lambda}$. This produces an element of $\tilde{\mathcal{F}}_{\lambda}$ contained in $\bigcap_{n<\omega}K_{n}$.
@@ 3.2 Filter Generation at Limit Stages
Let $\lambda$ be a limit ordinal and assume inductively that each $\mathcal{F}_{\beta}$ for $\beta<\lambda$ is a normal, $\omega_{1}$-complete filter on $\mathcal{G}_{\beta}$. We define the limit filter $\tilde{\mathcal{F}}_{\lambda}$ uniformly from the head pullbacks $\rho_{\beta,\lambda}^{-1}(H)$ (Definition 3.9). At $\mathrm{cf}(\lambda)\geq\omega_{1}$ we take the smallest normal filter containing these generators, and $\omega_{1}$-completeness follows by stage-bounding (Remark 3.5). At $\mathrm{cf}(\lambda)=\omega$ we instead take the smallest normal $\omega_{1}$-complete filter extending the same generator family.
Remark 3.6 (Head pullbacks of subgroups).
For $\beta<\lambda$ and $H\leq\mathcal{G}_{\beta}$, we write
$\rho_{\beta,\lambda}^{-1}(H)=\{g\in\mathcal{G}_{\lambda}:\rho_{\beta,\lambda}(g)\in H\}.$
When pulling back filters along homomorphisms we use the notation $\pi^{*}\mathcal{F}$ from Lemma 3.7.
@@
Definition 3.9 (Limit filter).
Let $\lambda$ be a limit ordinal and assume we have defined $(\mathbb{P}_{\beta},\mathcal{G}_{\beta},\mathcal{F}_{\beta})$ for all $\beta<\lambda$, with coherent restriction homomorphisms $\rho_{\beta,\lambda}:\mathcal{G}_{\lambda}\to\mathcal{G}_{\beta}$.
Define the head-generator family at $\lambda$ by
$\mathcal{B}_{\lambda}\ :=\ \bigl\{\ \rho_{\beta,\lambda}^{-1}(H)\ :\ \beta<\lambda\ \&\ H\in\mathcal{F}_{\beta}\ \bigr\}.$
@@
    If $\mathrm{cf}(\lambda)\geq\omega_{1}$, let $\tilde{\mathcal{F}}_{\lambda}$ be the normal filter on $\mathcal{G}_{\lambda}$ generated by $\mathcal{B}_{\lambda}$, equivalently:
@@
    If $\mathrm{cf}(\lambda)=\omega$, let $\tilde{\mathcal{F}}_{\lambda}$ be the normal $\omega_{1}$–complete filter on $\mathcal{G}_{\lambda}$ generated by $\mathcal{B}_{\lambda}$, equivalently:
@@
The $\omega_{1}$–completeness argument splits by cofinality. If $\mathrm{cf}(\lambda)\geq\omega_{1}$, stage-bounding reduces countable intersections in $\tilde{\mathcal{F}}_{\lambda}$ to a single stage filter (Remark 3.5). If $\mathrm{cf}(\lambda)=\omega$, $\omega_{1}$–completeness is built into the definition of $\tilde{\mathcal{F}}_{\lambda}$ as the smallest normal $\omega_{1}$–complete filter containing the head generators (Definition 3.9).
@@
When $\mathrm{cf}(\lambda)=\omega$, the stage-bounding lemma fails. For example, if $\lambda=\omega$, the set $\{n:n<\omega\}$ is countable but $\sup_{n}n=\omega=\lambda$. Similarly, if $\lambda=\aleph_{\omega}$, the sequence $\{\aleph_{n}:n<\omega\}$ is countable and cofinal. This is precisely why, at limits of countable cofinality, $\omega_{1}$–completeness must be built into the limit-filter definition (Definition 3.9(2)), rather than proved by stage-bounding.
Theorem 3.13 (Limit filter properties).
Let $\lambda$ be a limit stage and let $\tilde{\mathcal{F}}_{\lambda}$ be as in Definition 3.9. Assume inductively that each $\mathcal{F}_{\beta}$ for $\beta<\lambda$ is a normal $\omega_{1}$–complete filter on $\mathcal{G}_{\beta}$. Then:
@@
    $\tilde{\mathcal{F}}_{\lambda}$ is a normal filter on $\mathcal{G}_{\lambda}$ and contains every head pullback $\rho_{\beta,\lambda}^{-1}(H)$ for $\beta<\lambda$ and $H\in\mathcal{F}_{\beta}$.
@@
By Definition 3.9, $\tilde{\mathcal{F}}_{\lambda}$ is (by construction) the smallest normal filter (or normal $\omega_{1}$–complete filter) containing the generator family $\mathcal{B}_{\lambda}$, so it is a normal filter and contains each $\rho_{\beta,\lambda}^{-1}(H)$.
For $\omega_{1}$–completeness: if $\mathrm{cf}(\lambda)=\omega$, closure under countable intersections is built into the defining criterion in Definition 3.9(2) by interleaving witnesses. If $\mathrm{cf}(\lambda)\geq\omega_{1}$, $\omega_{1}$–completeness follows from stage-bounding as proved in Remark 3.5. ∎